A train is scheduled to travel with a velocity of 72 kmh-1 in a certain limit of its route. Due to renovation the



train had to travel 1km of the above limit with a velocity of 9 kmh-1.If the train travelled 1575m while



reducing velocity from 72 kmh-1to 9 kmh-1 and 3150m while increasing velocity from 9 kmh-1 to 72 km h-1



, how late will the train be?

1
Expert's answer
2022-02-06T14:32:42-0500

"\Delta t=t_2-t_1=\frac{2s_1}{v_2+v_1}+\frac{s_2}{v_1}+\frac{2s_3}{v_2+v_1}-\frac{s_1+s_2+s_3}{v_2}=8.9" min.